# Simultaneous Drawing of Layered Trees

### 12 Jan 2024 v3

@@ 2 Preliminaries
We further require that the total order $<_{1}$ of the layer $1$ (i.e., the order of all leaves) is given as part of the input, with the additional restriction that $<_{1}$ induces a planar embedding $\mathcal{E}_{i}$ with respect to each individual tree $T_{i}$, that is, there exists an ordering of the (non-leaf) vertices of $T_{i}$ such that no two edges of $T_{i}$ cross, see Fig. 1 for an illustration. Since the leaves of each $T_{i}$ are all on layer $1$, the embedding $\mathcal{E}_{i}$ is unique and implies a total ordering $<_{j}^{i}$ of the vertices of $T_{i}$ on every layer $j\in\{2,\dots,\ell\}$. Therefore, we henceforth assume that $V_{j}(T_{i})$ appears in the corresponding vertex order $<_{j}^{i}$, and if we combine all $<_{j}^{i}$ for $i\in\{1,\dots,k\}$, we obtain a partial order, which we call $\prec_{j}$.
@@ 3 Two Trees on Arbitrarily Many Layers
Consider some layer $j\in\{2,\dots,\ell\}$ and index the vertices in $V_{j}(T_{1})$ according $\mathcal{E}_{1}$ from left to right by $1,\dots,n_{1|j}$. In a complete drawing, we define, for a vertex $v$ of $T_{2}$, its position $p$ on layer $j$ with respect to the index of the closest vertex of $T_{1}$ to the left of $v$. If there is no such vertex, we set $p=0$. Let $C_{v}=\{c_{1},c_{2},\dots,c_{\deg(v)-1}\}$ be the ordered set of children of $v$ in $T_{2}$, which lie on layer $j-1$.
@@
A key observation is that for a position $p$ on layer $j$, there is precisely one ideal position $p^{\star}$ on layer $j-1$ such that $\operatorname{\chi}_{j-1}(p^{\star},p)=0$ and for two positions $p,q$ with $p<q$ on layer $j$, the ideal positions $p^{\star},q^{\star}$ on layer $j-1$ appear in the same order, i.e., $p^{\star}<q^{\star}$. (Imagine going down the gap of $\mathcal{E}_{1}$ where $p$ is located as illustrated in Fig. 1.) In Claim 1, we formalize another observation regarding ideal positions. Essentially, the claim says the further the endpoints of an edge are away from a pair of position and ideal position, the more crossings occur. For simplicity, we assume henceforth that each of the functions $o$ and $\operatorname{\chi}_{j}$ returns $\infty$ for parameters outside of its domain.
@@
Consider an edge $(u,v)$ of $T_{2}$ with its endpoints being placed at $p^{\star}$ and $p$. We know that $\operatorname{\chi}_{j-1}(p^{\star},p)=0$. Now, for every position that we move $u$ (resp. $v$) to the left or right of $p^{\star}$ (of $p$), we change sides with a vertex $w$ of $T_{1}$. Because $w$ has exactly (at least) one incident edge going upwards (downwards) to its parent (a child) that we have not crossed before, the number of crossings increases by exactly (at least) 1. ∎
@@
For a vertex $v\in V_{j}(T_{2})$ on a layer $j$, we define $\operatorname{P_{\textsf{opt}}}(v)$ as the set of every position $p$ where $o[v,p]$ is minimum. We analyze the properties of $\operatorname{P_{\textsf{opt}}}$ in Claim 2.
@@
For every layer $j\in\{2,\dots,L(r_{2})\}$, let $v_{1},v_{2},\dots,v_{n_{2|j}}$ be the vertices in $V_{j}(T_{2})$ in the order of $\mathcal{E}_{2}$. It holds that $\min\operatorname{P_{\textsf{opt}}}(v_{1})\leq\min\operatorname{P_{\textsf{opt}}}(v_{2})\leq\ldots\leq\min\operatorname{P_{\textsf{opt}}}(v_{n_{2|j}})$ and $\max\operatorname{P_{\textsf{opt}}}(v_{1})\leq\ldots\leq\max\operatorname{P_{\textsf{opt}}}(v_{n_{2|j}})$. Further, for every $v\in V_{j}(T_{2})$, $\operatorname{P_{\textsf{opt}}}(v)$ is an interval of natural numbers and, for any $x\in\mathbb{N}_{0}$, $o[v,\min\operatorname{P_{\textsf{opt}}}(v)-(x+1)]>o[v,\min\operatorname{P_{\textsf{opt}}}(v)-x]\geq x$ and $o[v,\max\operatorname{P_{\textsf{opt}}}(v)+(x+1)]>o[v,\max\operatorname{P_{\textsf{opt}}}(v)+x]\geq x$.
@@
We show this claim by induction over the layers $j=2,3,\dots$ For $j=2$ and every $v\in V_{j}(T_{2})$, the children of $v$ have fixed positions and, therefore, $o[v,p]$ only depends on the number of crossings induced by the position $p\in\{0,\dots,n_{1|j}\}$; see Fig. 2 for an example. We next show that $\operatorname{P_{\textsf{opt}}}(v)$ is an interval. Observe that $o[v,p]=\sum_{i=1}^{|C_{v}|}\operatorname{\chi}_{1}(p_{c_{i}},p)$ is a sum of discrete functions (with variable $p$) where each admits its minimum value for one or two neighboring values of $p$ and apart from at most two values at or around this minimum, all of these functions increase or decrease by the same amount if we add or subtract 1 to $p$, which follows by the argument presented in the proof of Claim 1. (These functions here are weakly unimodal, i.e., they have a global minimum and they increase monotonously when moving away from that minimum.) Now to find the positions where that sum is minimum, we traverse the values of its domain: we start with $p=0$. If we increase $p$ by one, then all functions that have not yet reached its minimum decrease, while the functions that had already reached their minimum increase by the same amount. Hence, this sum is minimum in the interval of the domain that has the minima of the single crossing functions equally distributed on the left and on the right side. Furthermore, for each position further to the left or right, the sum increases by at least one. It remains to show that the minima and maxima of $\operatorname{P_{\textsf{opt}}}(v_{1}),\operatorname{P_{\textsf{opt}}}(v_{2}),\dots,\operatorname{P_{\textsf{opt}}}(v_{n_{2|j}})$ increase monotonously. Since the children of the vertices on layer $2$ (i.e., the leaves) are ordered, the minima and maxima of all crossing functions are ordered and so are the minima and maxima of the sums. Now consider $j>2$. Again $o[v,p]$ is a sum, but now we add, for every child $c$ of $v$ and a position $q$, $o[c,q]$ and $\operatorname{\chi}_{j-1}(q,p)$. The sum of minima is again a sum of unimodal functions with similar properties as before: the $\operatorname{\chi}_{j-1}(q,p)$ summands increase and decrease around their minimum as before, while the $o[c,q]$ summands increase and decrease before and after their minimum at least as much as a $\operatorname{\chi}_{j-1}(q,p)$ summand due to the induction hypothesis. (They behave like a weighted $\operatorname{\chi}_{j-1}(q,p)$ summand.) Hence, if we sum them up, we apply a weighted version of the previous argument to obtain the properties stated in the claim. In particular, the minima and maxima of $\operatorname{P_{\textsf{opt}}}(v_{1}),\operatorname{P_{\textsf{opt}}}(v_{2}),\dots,\operatorname{P_{\textsf{opt}}}(v_{n_{2|j}})$ increase monotonously since the minima and maxima of $\operatorname{P_{\textsf{opt}}}$ of the children on layer $j-1$ are ordered by the induction hypothesis. ∎
@@
For a vertex $u$ on a layer $j-1$, we define the natural position $\operatorname{p_{\textsf{nat}}}(u,p)$ of $u$ with respect to the position $p$ of its parent vertex on layer $j$ as
$\operatorname{p_{\textsf{nat}}}(u,p)=\begin{cases}p^{\star},&\textrm{ if }p^{\star}\in\operatorname{P_{\textsf{opt}}}(u)\\ \max\operatorname{P_{\textsf{opt}}}(u),&\textrm{ if }p^{\star}>\max\operatorname{P_{\textsf{opt}}}(u)\\ \min\operatorname{P_{\textsf{opt}}}(u),&\textrm{ if }p^{\star}<\min\operatorname{P_{\textsf{opt}}}(u).\end{cases}$
@@
For a vertex $v\in V_{j}(T_{2})$ on a layer $j$, let $c_{1},\dots,c_{|C_{v}|}$ be the children of $v$. For any position $p\in\{0,\dots,n_{1|j}\}$, it holds that $\operatorname{p_{\textsf{nat}}}(c_{1},p)\leq\ldots\leq\operatorname{p_{\textsf{nat}}}(c_{|C_{v}|},p)$ and $o[v,p]=\sum_{i=1}^{|C_{v}|}(o[c_{i},\operatorname{p_{\textsf{nat}}}(c_{i},p)]+\operatorname{\chi}_{j-1}(\operatorname{p_{\textsf{nat}}}(c_{i},p),p))$.
@@
We partition the children of $v$ into three groups: if for a child $c_{i}$ (where $i\in\{1,\dots,|C_{v}|\}$), $p^{\star}\in\operatorname{P_{\textsf{opt}}}(c_{i})$, we set $q_{i}=p^{\star}$. If for a child $c_{i}$, $p^{\star}>\max\operatorname{P_{\textsf{opt}}}(c_{i})$, we set $q_{i}=\max\operatorname{P_{\textsf{opt}}}(c_{i})$, and, symmetrically, if $p^{\star}<\min\operatorname{P_{\textsf{opt}}}(c_{i})$, we set $q_{i}=\min\operatorname{P_{\textsf{opt}}}(c_{i})$. By Claim 2, we observe that $q_{1}\leq\ldots\leq q_{|C_{v}|}$. Since $q_{i}=\operatorname{p_{\textsf{nat}}}(c_{i},p)$, this proves the first part of the claim. Now for the second part, if $o[v,p]\neq\sum_{i=1}^{|C_{v}|}(o[c_{i},q_{i}]+\operatorname{\chi}_{j-1}(q_{i},p))$, then for some $i$, $o[c_{i},q_{i}]+\operatorname{\chi}_{j-1}(q_{i},p)>\min_{q^{\prime}\in\{0,\dots,n_{1|j-1}\}}(o[c_{i},q^{\prime}]+\operatorname{\chi}_{j-1}(q^{\prime},p))$. Let $\hat{q}\in\{0,\dots,n_{1|j-1}\}$ be a position such that $o[c_{i},q_{i}]+\operatorname{\chi}_{j-1}(q_{i},p)>o[c_{i},\hat{q}]+\operatorname{\chi}_{j-1}(\hat{q},p)$. Since we know $o[c_{i},q_{i}]\leq o[c_{i},\hat{q}]$, it follows that $\operatorname{\chi}_{j-1}(q_{i},p)>\operatorname{\chi}_{j-1}(\hat{q},p)$. First note that $p^{\star}\notin\operatorname{P_{\textsf{opt}}}(c_{i})$ because otherwise $\operatorname{\chi}_{j-1}(q_{i},p)<\operatorname{\chi}_{j-1}(\hat{q},p)$. It follows that $q_{i}$ is the minimum or maximum position of $\operatorname{P_{\textsf{opt}}}(c_{i})$ – assume w.l.o.g. that $q_{i}=\max\operatorname{P_{\textsf{opt}}}(c_{i})$. Then, $q_{i}<\hat{q}$ (and hence $\hat{q}\notin\operatorname{P_{\textsf{opt}}}(c_{i})$) because otherwise again $\operatorname{\chi}_{j-1}(q_{i},p)<\operatorname{\chi}_{j-1}(\hat{q},p)$. We distinguish two cases. The first case is $q_{i}<\hat{q}\leq p^{\star}$. By Claim 1, we know that $\operatorname{\chi}_{j-1}(q_{i},p)-\operatorname{\chi}_{j-1}(\hat{q},p)=\hat{q}-q_{i}$. By Claim 2, we know that $o[c_{i},\hat{q}]-o[c_{i},q_{i}]\geq\hat{q}-q_{i}$. Hence, we have $o[c_{i},\hat{q}]-o[c_{i},q_{i}]\geq\operatorname{\chi}_{j-1}(q_{i},p)-\operatorname{\chi}_{j-1}(\hat{q},p)$, which we can reformulate as $o[c_{i},q_{i}]+\operatorname{\chi}_{j-1}(q_{i},p)\leq o[c_{i},\hat{q}]+\operatorname{\chi}_{j-1}(\hat{q},p)$, which contradicts our initial assumption. The second case is $q_{i}<p^{\star}<\hat{q}$. Now we have $\operatorname{\chi}_{j-1}(q_{i},p)=p^{\star}-q_{i}$ and $\operatorname{\chi}_{j-1}(\hat{q},p)=\hat{q}-p^{\star}$. If we add up these two equations, we get $\operatorname{\chi}_{j-1}(q_{i},p)+\operatorname{\chi}_{j-1}(\hat{q},p)=\hat{q}-q_{i}$. As we still have $o[c_{i},\hat{q}]-o[c_{i},q_{i}]\geq\hat{q}-q_{i}$, we get $o[c_{i},q_{i}]+\operatorname{\chi}_{j-1}(q_{i},p)\leq o[c_{i},\hat{q}]-\operatorname{\chi}_{j-1}(\hat{q},p)$, which, of course, also contradicts our initial assumption. ∎
@@
For a vertex $v\in V_{j}(T_{2})$ on a layer $j$ and a position $p\in\{0,\dots,n_{1|j}\}$, the dynamic program selects $\operatorname{p_{\textsf{nat}}}(c_{1},p),\dots,\operatorname{p_{\textsf{nat}}}(c_{|C_{v}|},p)$ as the positions of $v$’s children $c_{1},\dots,c_{|C_{v}|}$ on layer $j-1$.
@@
Recall that, for any $i\in\{1,\dots,|C_{v}|\}$, if there is more than one position for $c_{i}$ resulting in a minimum value of $o[v,p]$, the position of $q$ with the maximum value of $\operatorname{\chi}_{j-1}(q,p)$ is used as a tie-breaker rule. If $\operatorname{p_{\textsf{nat}}}(c_{i},p)=p^{\star}$, then $p^{\star}$ is the only position of $c_{i}$ that can lead to a minimum value of $o[v,p]$ and our claim is true. Now, due to symmetry, we assume w.l.o.g. that $\operatorname{p_{\textsf{nat}}}(c_{i},p)=\max\operatorname{P_{\textsf{opt}}}(c_{i})$. Let $p^{\prime}\neq\operatorname{p_{\textsf{nat}}}(c_{i},p)$ be a position of $c_{i}$ yielding a minimum value of $o[v,p]$. The position $p^{\prime}$ cannot lie within $\operatorname{P_{\textsf{opt}}}(c_{i})$ by Claim 1 since this would result in a larger number of crossings, while $o[c_{i},p^{\prime}]=o[c_{i},\operatorname{p_{\textsf{nat}}}(c_{i},p)]$. Hence, $\operatorname{p_{\textsf{nat}}}(c_{i},p)<p^{\prime}$. By Claim 2, $o[c_{i},p^{\prime}]-o[c_{i},\operatorname{p_{\textsf{nat}}}(c_{i},p)]\geq p^{\prime}-\operatorname{p_{\textsf{nat}}}(c_{i},p)$. This means, that, for each position further to the right of $\operatorname{p_{\textsf{nat}}}(c_{i},p)$, the value of the dynamic program for $c_{i}$ increases by at least one, while the number of crossings according to the function $\operatorname{\chi}_{j-1}$ increases by exactly one (see Claim 1). Thus, $\operatorname{p_{\textsf{nat}}}(c_{i},p)$ is one (of possibly several) position(s) of $c_{i}$ admitting a minimum value of $o[v,p]$. If $p^{\prime}$ also admits a minimum value of $o[v,p]$, but $o[c_{i},p^{\prime}]>o[c_{i},\operatorname{p_{\textsf{nat}}}(c_{i},p)]$, it follows that $\operatorname{\chi}_{j-1}(p^{\prime},p)<\operatorname{\chi}_{j-1}(\operatorname{p_{\textsf{nat}}}(c_{i},p),p)$. Hence, due to the tie-breaker rule, our algorithm would have selected $\operatorname{p_{\textsf{nat}}}(c_{i},p)$ instead of $p^{\prime}$. ∎

### 28 Feb 2024 v4

@@ 2 Preliminaries
We further require that the total order $<_{1}$ of layer $1$ (i.e., the order of all leaves) is given as part of the input, with the additional restriction that $<_{1}$ induces a planar embedding $\mathcal{E}_{i}$ with respect to each individual tree $T_{i}$, that is, there exists an ordering of the (non-leaf) vertices of $T_{i}$ such that no two edges of $T_{i}$ cross, see Fig. 1 for an illustration. Since the leaves of each $T_{i}$ are all on layer $1$, the embedding $\mathcal{E}_{i}$ is unique and implies a total ordering $<_{j}^{i}$ of the vertices of $T_{i}$ on every layer $j\in\{2,\dots,\ell\}$. Therefore, we henceforth assume that $V_{j}(T_{i})$ appears in the corresponding vertex order $<_{j}^{i}$, and if we combine all $<_{j}^{i}$ for $i\in\{1,\dots,k\}$, we obtain a partial order, which we call $\prec_{j}$.
@@ 3 Two Trees on Arbitrarily Many Layers
Consider some layer $j\in\{2,\dots,\ell\}$ and index the vertices in $V_{j}(T_{1})$ according $\mathcal{E}_{1}$ from left to right by $1,\dots,n_{1|j}$. In a complete drawing, we define, for a vertex $v$ of $T_{2}$, its position $p$ on layer $j$ with respect to the index of the closest vertex of $T_{1}$ to the left of $v$. If there is no such vertex, we set $p=0$. Let $C_{v}=\{c_{1},c_{2},\dots,c_{\mathrm{indeg}(v)}\}$ be the ordered set of children of $v$ in $T_{2}$, which lie on layer $j-1$, where $\mathrm{indeg}(v)$ is the in-degree of $v$.
@@
A key observation is that for a position $p$ on layer $j$, there is precisely one ideal position $p^{\star}$ on layer $j-1$ such that $\operatorname{\chi}_{j-1}(p^{\star},p)=0$ and for two positions $p,q$ with $p<q$ on layer $j$, the ideal positions $p^{\star},q^{\star}$ on layer $j-1$ appear in the same order, i.e., $p^{\star}<q^{\star}$. (Imagine going down the gap of $\mathcal{E}_{1}$ where $p$ is located as illustrated in Fig. 1.) In Claim 1, we formalize another observation regarding ideal positions. Essentially, the claim says that the further the endpoints of an edge are away from a pair of position and ideal position, the more crossings occur. For simplicity, we assume henceforth that each of the functions $o$ and $\operatorname{\chi}_{j}$ returns $\infty$ for parameters outside of its domain.
@@
Consider an edge $(u,v)$ of $T_{2}$ with its endpoints being placed at $p^{\star}$ and $p$. We know that $\operatorname{\chi}_{j-1}(p^{\star},p)=0$. Now, for every position that we move $u$ ($v$, resp.) to the left or right of $p^{\star}$ (of $p$), we change sides with a vertex $w$ of $T_{1}$. Because $w$ has exactly (at least) one incident edge going upwards (downwards) to its parent (a child) that we have not crossed before, the number of crossings increases by exactly (at least) 1. ∎
@@
For a vertex $v\in V_{j}(T_{2})$ on a layer $j$, we define $\operatorname{P_{\mathsf{opt}}}(v)$ as the set of every position $p$ where $o[v,p]$ is minimum. We analyze the properties of $\operatorname{P_{\mathsf{opt}}}$ in Claim 2.
@@
For every layer $j\in\{2,\dots,L(r_{2})\}$, let $v_{1},v_{2},\dots,v_{n_{2|j}}$ be the vertices in $V_{j}(T_{2})$ in the order of $\mathcal{E}_{2}$. It holds that $\min\operatorname{P_{\mathsf{opt}}}(v_{1})\leq\min\operatorname{P_{\mathsf{opt}}}(v_{2})\leq\ldots\leq\min\operatorname{P_{\mathsf{opt}}}(v_{n_{2|j}})$ and $\max\operatorname{P_{\mathsf{opt}}}(v_{1})\leq\ldots\leq\max\operatorname{P_{\mathsf{opt}}}(v_{n_{2|j}})$. Further, for every $v\in V_{j}(T_{2})$, $\operatorname{P_{\mathsf{opt}}}(v)$ is an interval of natural numbers and, for any $x\in\mathbb{N}_{0}$, $o[v,\min\operatorname{P_{\mathsf{opt}}}(v)-(x+1)]>o[v,\min\operatorname{P_{\mathsf{opt}}}(v)-x]\geq x$ and $o[v,\max\operatorname{P_{\mathsf{opt}}}(v)+(x+1)]>o[v,\max\operatorname{P_{\mathsf{opt}}}(v)+x]\geq x$.
@@
We show this claim by induction over the layers $j=2,3,\dots$ For $j=2$ and every $v\in V_{j}(T_{2})$, the children of $v$ have fixed positions and, therefore, $o[v,p]$ only depends on the number of crossings induced by the position $p\in\{0,\dots,n_{1|j}\}$; see Fig. 2 for an example. We next show that $\operatorname{P_{\mathsf{opt}}}(v)$ is an interval. Observe that $o[v,p]=\sum_{i=1}^{|C_{v}|}\operatorname{\chi}_{1}(p_{c_{i}},p)$ is a sum of discrete functions (with variable $p$) where each admits its minimum value for one or two neighboring values of $p$ and apart from at most two values at or around this minimum, all of these functions increase or decrease by the same amount if we add or subtract 1 to $p$, which follows by the argument presented in the proof of Claim 1. (These functions here are weakly unimodal, i.e., they have a global minimum and they increase monotonously when moving away from that minimum.) Now to find the positions where that sum is minimum, we traverse the values of its domain: we start with $p=0$. If we increase $p$ by one, then all functions that have not yet reached its minimum decrease, while the functions that had already reached their minimum increase by the same amount. Hence, this sum is minimum in the interval of the domain that has the minima of the single crossing functions equally distributed on the left and on the right side. Furthermore, for each position further to the left or right, the sum increases by at least one. It remains to show that the minima and maxima of $\operatorname{P_{\mathsf{opt}}}(v_{1}),\operatorname{P_{\mathsf{opt}}}(v_{2}),\dots,\operatorname{P_{\mathsf{opt}}}(v_{n_{2|j}})$ increase monotonously. Since the children of the vertices on layer $2$ (i.e., the leaves) are ordered, the minima and maxima of all crossing functions are ordered and so are the minima and maxima of the sums. Now consider $j>2$. Again $o[v,p]$ is a sum, but now we add, for every child $c$ of $v$ and a position $q$, $o[c,q]$ and $\operatorname{\chi}_{j-1}(q,p)$. The sum of minima is again a sum of unimodal functions with similar properties as before: the $\operatorname{\chi}_{j-1}(q,p)$ summands increase and decrease around their minimum as before, while the $o[c,q]$ summands increase and decrease before and after their minimum at least as much as a $\operatorname{\chi}_{j-1}(q,p)$ summand due to the induction hypothesis. (They behave like a weighted $\operatorname{\chi}_{j-1}(q,p)$ summand.) Hence, if we sum them up, we apply a weighted version of the previous argument to obtain the properties stated in the claim. In particular, the minima and maxima of $\operatorname{P_{\mathsf{opt}}}(v_{1}),\operatorname{P_{\mathsf{opt}}}(v_{2}),\dots,\operatorname{P_{\mathsf{opt}}}(v_{n_{2|j}})$ increase monotonously since the minima and maxima of $\operatorname{P_{\mathsf{opt}}}$ of the children on layer $j-1$ are ordered by the induction hypothesis. ∎
@@
For a vertex $u$ on a layer $j-1$, we define the natural position $\operatorname{p_{\mathsf{nat}}}(u,p)$ of $u$ with respect to the position $p$ of its parent vertex on layer $j$ as
$\operatorname{p_{\mathsf{nat}}}(u,p)=\begin{cases}p^{\star},&\textrm{ if }p^{\star}\in\operatorname{P_{\mathsf{opt}}}(u)\\ \max\operatorname{P_{\mathsf{opt}}}(u),&\textrm{ if }p^{\star}>\max\operatorname{P_{\mathsf{opt}}}(u)\\ \min\operatorname{P_{\mathsf{opt}}}(u),&\textrm{ if }p^{\star}<\min\operatorname{P_{\mathsf{opt}}}(u).\end{cases}$
@@
For a vertex $v\in V_{j}(T_{2})$ on a layer $j$, let $c_{1},\dots,c_{|C_{v}|}$ be the children of $v$. For any position $p\in\{0,\dots,n_{1|j}\}$, it holds that $\operatorname{p_{\mathsf{nat}}}(c_{1},p)\leq\ldots\leq\operatorname{p_{\mathsf{nat}}}(c_{|C_{v}|},p)$ and $o[v,p]=\sum_{i=1}^{|C_{v}|}(o[c_{i},\operatorname{p_{\mathsf{nat}}}(c_{i},p)]+\operatorname{\chi}_{j-1}(\operatorname{p_{\mathsf{nat}}}(c_{i},p),p))$.
@@
We partition the children of $v$ into three groups: if for a child $c_{i}$ (where $i\in\{1,\dots,|C_{v}|\}$), $p^{\star}\in\operatorname{P_{\mathsf{opt}}}(c_{i})$, we set $q_{i}=p^{\star}$. If for a child $c_{i}$, $p^{\star}>\max\operatorname{P_{\mathsf{opt}}}(c_{i})$, we set $q_{i}=\max\operatorname{P_{\mathsf{opt}}}(c_{i})$, and, symmetrically, if $p^{\star}<\min\operatorname{P_{\mathsf{opt}}}(c_{i})$, we set $q_{i}=\min\operatorname{P_{\mathsf{opt}}}(c_{i})$. By Claim 2, we observe that $q_{1}\leq\ldots\leq q_{|C_{v}|}$. Since $q_{i}=\operatorname{p_{\mathsf{nat}}}(c_{i},p)$, this proves the first part of the claim. Now for the second part, if $o[v,p]\neq\sum_{i=1}^{|C_{v}|}(o[c_{i},q_{i}]+\operatorname{\chi}_{j-1}(q_{i},p))$, then for some $i$, $o[c_{i},q_{i}]+\operatorname{\chi}_{j-1}(q_{i},p)>\min_{q^{\prime}\in\{0,\dots,n_{1|j-1}\}}(o[c_{i},q^{\prime}]+\operatorname{\chi}_{j-1}(q^{\prime},p))$. Let $\hat{q}\in\{0,\dots,n_{1|j-1}\}$ be a position such that $o[c_{i},q_{i}]+\operatorname{\chi}_{j-1}(q_{i},p)>o[c_{i},\hat{q}]+\operatorname{\chi}_{j-1}(\hat{q},p)$. Since we know $o[c_{i},q_{i}]\leq o[c_{i},\hat{q}]$, it follows that $\operatorname{\chi}_{j-1}(q_{i},p)>\operatorname{\chi}_{j-1}(\hat{q},p)$. First note that $p^{\star}\notin\operatorname{P_{\mathsf{opt}}}(c_{i})$ because otherwise $\operatorname{\chi}_{j-1}(q_{i},p)<\operatorname{\chi}_{j-1}(\hat{q},p)$. It follows that $q_{i}$ is the minimum or maximum position of $\operatorname{P_{\mathsf{opt}}}(c_{i})$ – assume w.l.o.g. that $q_{i}=\max\operatorname{P_{\mathsf{opt}}}(c_{i})$. Then, $q_{i}<\hat{q}$ (and hence $\hat{q}\notin\operatorname{P_{\mathsf{opt}}}(c_{i})$) because otherwise again $\operatorname{\chi}_{j-1}(q_{i},p)<\operatorname{\chi}_{j-1}(\hat{q},p)$. We distinguish two cases. The first case is $q_{i}<\hat{q}\leq p^{\star}$. By Claim 1, we know that $\operatorname{\chi}_{j-1}(q_{i},p)-\operatorname{\chi}_{j-1}(\hat{q},p)=\hat{q}-q_{i}$. By Claim 2, we know that $o[c_{i},\hat{q}]-o[c_{i},q_{i}]\geq\hat{q}-q_{i}$. Hence, we have $o[c_{i},\hat{q}]-o[c_{i},q_{i}]\geq\operatorname{\chi}_{j-1}(q_{i},p)-\operatorname{\chi}_{j-1}(\hat{q},p)$, which we can reformulate as $o[c_{i},q_{i}]+\operatorname{\chi}_{j-1}(q_{i},p)\leq o[c_{i},\hat{q}]+\operatorname{\chi}_{j-1}(\hat{q},p)$, which contradicts our initial assumption. The second case is $q_{i}<p^{\star}<\hat{q}$. Now we have $\operatorname{\chi}_{j-1}(q_{i},p)=p^{\star}-q_{i}$ and $\operatorname{\chi}_{j-1}(\hat{q},p)=\hat{q}-p^{\star}$. If we add up these two equations, we get $\operatorname{\chi}_{j-1}(q_{i},p)+\operatorname{\chi}_{j-1}(\hat{q},p)=\hat{q}-q_{i}$. As we still have $o[c_{i},\hat{q}]-o[c_{i},q_{i}]\geq\hat{q}-q_{i}$, we get $o[c_{i},q_{i}]+\operatorname{\chi}_{j-1}(q_{i},p)\leq o[c_{i},\hat{q}]-\operatorname{\chi}_{j-1}(\hat{q},p)$, which, of course, also contradicts our initial assumption. ∎
@@
For a vertex $v\in V_{j}(T_{2})$ on a layer $j$ and a position $p\in\{0,\dots,n_{1|j}\}$, the dynamic program selects $\operatorname{p_{\mathsf{nat}}}(c_{1},p),\dots,\operatorname{p_{\mathsf{nat}}}(c_{|C_{v}|},p)$ as the positions of $v$’s children $c_{1},\dots,c_{|C_{v}|}$ on layer $j-1$.
@@
Recall that, for any $i\in\{1,\dots,|C_{v}|\}$, if there is more than one position for $c_{i}$ resulting in a minimum value of $o[v,p]$, the position of $q$ with the maximum value of $\operatorname{\chi}_{j-1}(q,p)$ is used as a tie-breaker rule. If $\operatorname{p_{\mathsf{nat}}}(c_{i},p)=p^{\star}$, then $p^{\star}$ is the only position of $c_{i}$ that can lead to a minimum value of $o[v,p]$ and our claim is true. Now, due to symmetry, we assume w.l.o.g. that $\operatorname{p_{\mathsf{nat}}}(c_{i},p)=\max\operatorname{P_{\mathsf{opt}}}(c_{i})$. Let $p^{\prime}\neq\operatorname{p_{\mathsf{nat}}}(c_{i},p)$ be a position of $c_{i}$ yielding a minimum value of $o[v,p]$. The position $p^{\prime}$ cannot lie within $\operatorname{P_{\mathsf{opt}}}(c_{i})$ by Claim 1 since this would result in a larger number of crossings, while $o[c_{i},p^{\prime}]=o[c_{i},\operatorname{p_{\mathsf{nat}}}(c_{i},p)]$. Hence, $\operatorname{p_{\mathsf{nat}}}(c_{i},p)<p^{\prime}$. By Claim 2, $o[c_{i},p^{\prime}]-o[c_{i},\operatorname{p_{\mathsf{nat}}}(c_{i},p)]\geq p^{\prime}-\operatorname{p_{\mathsf{nat}}}(c_{i},p)$. This means, that, for each position further to the right of $\operatorname{p_{\mathsf{nat}}}(c_{i},p)$, the value of the dynamic program for $c_{i}$ increases by at least one, while the number of crossings according to the function $\operatorname{\chi}_{j-1}$ increases by exactly one (see Claim 1). Thus, $\operatorname{p_{\mathsf{nat}}}(c_{i},p)$ is one (of possibly several) position(s) of $c_{i}$ admitting a minimum value of $o[v,p]$. If $p^{\prime}$ also admits a minimum value of $o[v,p]$, but $o[c_{i},p^{\prime}]>o[c_{i},\operatorname{p_{\mathsf{nat}}}(c_{i},p)]$, it follows that $\operatorname{\chi}_{j-1}(p^{\prime},p)<\operatorname{\chi}_{j-1}(\operatorname{p_{\mathsf{nat}}}(c_{i},p),p)$. Hence, due to the tie-breaker rule, our algorithm would have selected $\operatorname{p_{\mathsf{nat}}}(c_{i},p)$ instead of $p^{\prime}$. ∎